# Scale-invariant attention

### v1

@@ 1 Introduction
One key challenge in modern LLMs is scaling up context length at inference time, while maintaining model performance. We approach this question of length generalisation by considering scale invariance. In particular, we are inspired by the “scale-invariant” statistics of natural images (Van der Schaaf & van Hateren, 1996). Scale invariance, for images, is actually highly intuitive, and means that there is structure at all spatial scales. For example, in an image there might be big features that are 100–1000 pixels across, and some small features that are only 1–10 pixels across. In natural images, features at both spatial scales are important: you cannot remove features at either scale without radically altering the image (Van der Schaaf & van Hateren, 1996).
@@
Total attention tells us the amount of attention a certain region receives. However, scale invariance of the total attention does not tell us about how the attention will be distributed among tokens in the region, i.e. whether the attention is spread out among many tokens or concentrated onto only a few tokens. As the region gets wider (e.g. from 10–100 tokens ago to 10,000–100,000 tokens ago), we might expect attention to spread out over more tokens simply due to the increased number of tokens. This “spreading out” of attention is possibly suboptimal for large contexts (Nakanishi, 2025), and that instead, we should focus attention on only a few of the most relevant tokens. To measure these effects, we use the entropy, which roughly captures the logarithm of the number of tokens attended to.
@@
  • We implement scale-invariant attention in conjunction with $p$-RoPE (Barbero et al., 2024b). We show that our method, ‘scale-invariant $p$-RoPE’, exhibits improvements in validation loss both when doing long-context training, and when zero-shot generalising to longer contexts. Our method also matches the performance of the best alternatives in an out-of-distribution long context ‘needle-in-a-haystack’ task.
@@ 2 Related work
Handling long sequences effectively in Transformer-based models remains a significant challenge and is of high interest to the deep learning community (Bai et al., 2024; Ye et al., 2024; Jin et al., 2024; Beltagy et al., 2020; Ding et al., 2023; Munkhdalai et al., 2024; Bulatov et al., 2024; Liu et al., 2023a; Barbero et al., 2024a; Hu et al., 2024). Below, we discuss several strategies in this literature relating to our work.
Long contexts via the positional encoding: Methods like ALiBi (Press et al., 2021) introduce a static, causal bias directly into the attention logits. ALiBi endows the model with an inductive bias towards recent tokens (Kazemnejad et al., 2023), thus helping with longer contexts. Rotary Position Embeddings (RoPE) (Su et al., 2024) have emerged as the most popular position encoding scheme, encoding relative positions. RoPE does not generalise to longer sequences out of the box (Peng et al., 2023), leading to subsequent research improving RoPE’s length extrapolation capabilities (Zhu et al., 2023; Wang et al., 2024). Many achieve this by modifying RoPE’s frequency spectrum; for example, Positional Interpolation (PI) (Chen et al., 2023), NTK-aware scaling (bloc97, 2023), YaRN (Peng et al., 2023), and by simply increasing the RoPE base $\theta$ parameter (Grattafiori et al., 2024; Gemma Team, 2025; Barbero et al., 2024b; Liu et al., 2023b).
Entropy Control: Beyond positional information, the properties of the attention distribution itself are crucial, especially in long contexts where attention scores can “smear”/“spread out” across many tokens. Scalable-Softmax (SSMax), also known as the ‘LogN trick’ (Nakanishi, 2025; Chiang & Cholak, 2022; Jianlin, 2021; Bai et al., 2023; Llama 4 Team, 2025), addresses this by multiplying the attention logits by $s\log N$, where $N$ is the context length and $s$ is a learned scale parameter. This multiplier has the effect of sharpening/focusing the attention distribution. Li et al. (2025) adopt a similar approach: controlling the entropy of the attention distribution for better length generalization.
@@
Efficient Long-Context Training and Inference: A key observation of Xiong et al. (2023) is that continual pretraining on long contexts, after initial pretraining on shorter sequences, is often sufficient and much more computationally efficient. Thus, a simple strategy for improving long-context performance, given sufficient computational budget, is continual pretraining on longer contexts. This has been adopted widely (Grattafiori et al., 2024; Gao et al., 2024; Lieber et al., 2024; Yang et al., 2024; Cohere Team, 2025; Llama 4 Team, 2025; Liu et al., 2024a). Of course, this strategy considerably increases the complexity and memory cost of the pretraining pipeline, making approaches like ours that can zero-shot generalise very valuable. Furthermore, one might expect long-context training to be far easier (e.g. requiring fewer long-context training steps for optimal performance) for approaches that already have good long-context performance due to zero-shot generalisation.
For inference on sequences exceeding the trained context length, alternative strategies bypass applying dense attention over the entire context, allowing for ‘infinite attention’ (Munkhdalai et al., 2024; Liu et al., 2024b; Martins et al., 2021; Chen et al., 2025; Ding et al., 2023). These include maintaining a fixed-size attention window (Beltagy et al., 2020), retrieving relevant context tokens using heuristics like top-K proximity (Han et al., 2023), or vector similarity search akin to episodic memory systems (Fountas et al., 2024; Xiao et al., 2024). While effective, these approaches operate at a higher level, managing context rather than modifying the core attention mechanism’s ability to process it directly, which is the focus of our work.
@@ 3 Methods
Following FlexAttention (Dong et al., 2024), we define the attention “score” as the dot product of a query $q$ and keys $K$,
@@ 3.1 Formal definitions of scale-invariant total attention and attention sparsity
The “$\Theta(1)$” is big-$\Theta$ notation, and means there exist constants $c_{1},c_{2}>0$ and $t_{0}$ such that for all $t>t_{0}$, $c_{1}\leq\operatorname{E}\left[Z_{t}^{t\Delta}\right]\leq c_{2}$.
@@
Figure 2: Expected entropy of scale-invariant attention at different scales is sub-logarithmic. Here, we sample sequences of independent standard Gaussian logits, and apply the scale-invariant attention transformation. We estimate the expected entropy in ranges $[t,t\Delta)$, where the size of the range is controlled by $t$ (x-axis) and $\Delta$ (line color). We see that this expected entropy measure scales sub-logarithmically (left), and with the right plot suggesting a $\sim\sqrt{\log(t)}$ scaling. The dashed lines show a best linear fit.
@@ 3.2 What characteristics of the logits are required for scale-invariant attention?
then we have scale-invariant unnormalised attention sparsity (Def. 3.2). To construct an attention mechanism that satisfies Eq. (12) and Eq. (13), we consider a simplified setting in which the logits are marginally Gaussian and arise from taking IID Gaussian “base logits”, $\bar{L}_{t}\sim\mathcal{N}\left(0,1\right)$, and transforming them by multiplying by $a_{t}$ and adding a bias, $m_{t}$,
@@
For this solution to be valid, we only require $\beta\geq\alpha\log\alpha$ since $\log(t/\tau+1)\geq 0$ when $t\geq 0$. We formally summarise the above results in Theorem 16 (proof in Appendix F), which tells us that this approach does indeed give scale-invariant total attention and scale-invariant unnormalised attention sparsity. {restatable}theoremthmscaleinv Suppose attention logits $\{L_{t}\}$ are independent and marginally Gaussian with mean $m_{t}$ and standard deviation $a_{t}$ defined by Eq. (16). Assuming $\alpha,\beta,\tau>0$, $\beta\geq\alpha\log\alpha$, then we have scale-invariant total attention and scale-invariant unnormalised attention sparsity. We therefore propose to scale the logits in real attention using $a_{t}$ and $m_{t}$ defined in Eq. (16). As $t$ increases, the variance, $a_{t}^{2}$, increases as the logarithm of $t$, while the mean decreases as the logarithm of $t$.
@@ 4 Experiments
In this section, we compare scale-invariant attention with other dense attention methods including Dynamic NTK interpolation (RoPE+NTK) (bloc97, 2023), LogN scaling/SSMax (Nakanishi, 2025), $p$-RoPE (Barbero et al., 2024b), and ALiBi (Press et al., 2021). Our results show that our method, scale-invariant $p$-RoPE, has uniformly lower validation loss at a variety of training lengths (4k, 16k, 64k). Additionally, scale-invariant $p$-RoPE demonstrates stronger zero-shot long-context generalization (e.g. in ‘Train @4k/Val @16k’ and ‘Train @4k/Val @64k’ settings) versus all other methods. Finally, scale-invariant $p$-RoPE, along with LogN, saturated“needle-in-a-haystack” task.
We pretrained GPT-2-style models (Radford et al., 2019) (with QK-norm, ReLU2 activations, etc. (Jordan et al., 2024a)) from scratch on the FineWeb dataset (Penedo et al., 2024), using a fixed training data ordering and a 10M token validation set. We trained linear layers with Muon (Jordan et al., 2024b), and remaining parameters with Adam. We implemented scale-invariant attention using FlexAttention (Dong et al., 2024).
We trained two models: one with 162M parameters and another with 304M. The 162M one were trained for 4578 steps on 2.4B tokens over a range of context lengths (4k, 16k, 64k). The 304M parameter models were trained only on the best length-generalising methods, for 10.9k steps on 10B tokens, at 4k context length. We give further experimental details in Appendix G. For the 162M parameter model, we targeted a validation loss of 3.28, following Karpathy’s GPT-2 reproduction (Karpathy, 2024; Jordan et al., 2024a), shown in the figures as a horizontal grey dashed line.
@@
To assess whether long-context information retrieval capabilities suffered, we fine-tuned models on a needle-in-a-haystack task (note that fine-tuning on this task is unusual, but we found that prompting alone was not sufficient to perform this task, as bigger models/more pretraining would be required). Needle-in-a-haystack (Kamradt, 2023) measures a model’s ability to precisely retrieve specific details (needles) from a large body of text (the haystack). Our tasks constructs prompts by concatenating text samples from the C4 dataset (Roberts et al., 2019) and embedding “needles” of the form ‘The special magic <city> number is <7_digit_number>’. We insert three (rather than one) needles uniformly, at random, into each context for more signal per example, with each needle contributing separately to the overall accuracy. A successful retrieval requires the model to output both the city and the associated number correctly. We trained models on sequences of length 4k, and tested at 4k, 16k, and 64k.
@@
Table 1: Final mean validation losses (±$\pm 1$ standard error across 3 seeds) for different methods when training with 4k context length on a 162M parameter GPT-2-style model. The error bars are small due to consistent training data ordering, and a fixed validation set.
@@
RoPE 0.962 $\pm$ 0.003 $0$$\pm$ 0.000 $0$$\pm$ 0.000
$p$-RoPE 0.966 $\pm$ 0.001 0.250 $\pm$ 0.020 $0$$\pm$ 0.000
NoPE 0.964 $\pm$ 0.000 0.303 $\pm$ 0.239 $0$$\pm$ 0.000
RoPE+NTK 0.962 $\pm$ 0.001 0.217 $\pm$ 0.078 $0$$\pm$ 0.000
@@ 5 Limitations
In this work, we evaluated our methods using models with 162M and 304M parameters. Ideally, we would have pretrained models at the multi-billion parameter scale, in line with contemporary state-of-the-art LLMs. While compute resource constraints prevented this, our results, together with the natural theoretical approach, offer no indication that our conclusions would fail to generalise to larger, commercially deployed models.
@@ Appendix A Relating the Indicies in the Paper to Standard Indices
where $\lambda$ indexes the feature, $i$ indexes the query (i.e. the token we’re generating now) and $j$ indexes the key (i.e. the token we’re attending to). We fix $i$, and take $t=i-j$, so
@@ Appendix B $\operatorname{E}\left[X^{k}\exp(\alpha X)\right]$ where $X$ is Gaussian
[…]
@@ Appendix C Deriving $a_{t}$ and $m_{t}$ in the Logit Transformation
and $\text{and }a_{t}^{2}$ $=\frac{\beta}{\alpha}-m_{t}.$ (35)
@@ G.1 Pretraining from scratch
Our base model is a modded-nanogpt (Jordan et al., 2024a) variant, which is similar to GPT-2 (Radford et al., 2019), but has the following main differences from GPT-2: RMSNorm for layer normalization (applied before the attention and MLP blocks, as well as to the token embeddings and the final output layer), squared ReLU activations, and QK Normalization (RMSNorm on query and key projections). All models used a vocabulary size of 50304, with text tokenized with the GPT-2 tokenizer (Wolf et al., 2020). We implemented scale-invariant attention and ALiBi using FlexAttention (Dong et al., 2024).
@@
304M parameter model. For the larger model, we used the same settings as the smaller model, but with the following changes. The larger model had 16 layers, 1024 hidden dimension, and 8 heads. We trained the model for 10900 steps, processing approximately 10B tokens. We trained for 2 accumulation steps on 4 GPUs, with 28 sequences per batch. We scaled learning rates following $\mu$Param (Wortsman et al., 2023), which involves multiplying learning rates of the linear layers by $768/1024$, to adjust for changing the model width. Learning rates of other layers (embedding, unembedding, and LogN scalings if active) were not changed. We used a cosine learning rate schedule, with no warmup, and a minimum learning rate of 0 (at the end of training). All runs on the 304M parameter models were with 4096 training sequence length. We validated at 4k, 16k, 64k sequence lengths every 250 steps.
@@
Dataset. We used subsets of FineWeb (Penedo et al., 2024) for pretraining from scratch: a 10B token subset for the 162M model, and a 100B token subset for the 304M model (from which $\sim$10B tokens were used for its training). We keep the training data ordering fixed, and we keep the validation set identical across all runs to reduce variance.

### v2

@@ 1 Introduction
One key challenge in modern LLMs is scaling up context length at inference time, while maintaining model performance. We approach this question of length generalisation by considering scale invariance. In particular, we are inspired by the “scale-invariant” statistics of natural images (van1996modelling). Scale invariance, for images, is actually highly intuitive, and means that there is structure at all spatial scales. For example, in an image there might be big features that are 100–1000 pixels across, and some small features that are only 1–10 pixels across. In natural images, features at both spatial scales are important: you cannot remove features at either scale without radically altering the image (van1996modelling).
@@
Total attention tells us the amount of attention a certain region receives. However, scale invariance of the total attention does not tell us about how the attention will be distributed among tokens in the region, i.e. whether the attention is spread out among many tokens or concentrated onto only a few tokens. As the region gets wider (e.g. from 10–100 tokens ago to 10,000–100,000 tokens ago), we might expect attention to spread out over more tokens simply due to the increased number of tokens. This “spreading out” of attention is possibly suboptimal for large contexts (nakanishi2025scalable), and that instead, we should focus attention on only a few of the most relevant tokens. To measure these effects, we use the entropy, which roughly captures the logarithm of the number of tokens attended to.
@@
  • We implement scale-invariant attention in conjunction with $p$-RoPE (barbero2024round). We show that our method, ‘scale-invariant $p$-RoPE’, exhibits improvements in validation loss both when doing long-context training, and when zero-shot generalising to longer contexts. Our method also matches the performance of the best alternatives in an out-of-distribution long context ‘needle-in-a-haystack’ task.
@@ 2 Related work
Handling long sequences effectively in Transformer-based models remains a significant challenge and is of high interest to the deep learning community (bai2024longalign; ye2024mplug; jin2024llm; beltagy2020longformer; ding2023longnet; munkhdalai2024leave; bulatov2024beyond; liu2023ring; barbero2024transformers; hu2024longrecipe). Below, we discuss several strategies in this literature relating to our work.
Long contexts via the positional encoding: Methods like ALiBi (press2021train) introduce a static, causal bias directly into the attention logits. ALiBi endows the model with an inductive bias towards recent tokens (kazemnejad2023impact), thus helping with longer contexts. Rotary Position Embeddings (RoPE) (su2024roformer) have emerged as the most popular position encoding scheme, encoding relative positions. RoPE does not generalise to longer sequences out of the box (peng2023yarn), leading to subsequent research improving RoPE’s length extrapolation capabilities (zhu2023pose; wang2024resonance). Many achieve this by modifying RoPE’s frequency spectrum; for example, Positional Interpolation (PI) (chen2023extending), NTK-aware scaling (bloc97_2023), YaRN (peng2023yarn), and by simply increasing the RoPE base $\theta$ parameter (grattafiori2024llama; team2025gemma; barbero2024round; liu2023scaling).
Entropy Control: Beyond positional information, the properties of the attention distribution itself are crucial, especially in long contexts where attention scores can “smear”/“spread out” across many tokens. Scalable-Softmax (SSMax), also known as the ‘LogN trick’ (nakanishi2025scalable; chiang2022overcoming; kexuefm-8823; bai2023qwen; llama4team2025), addresses this by multiplying the attention logits by $s\log N$, where $N$ is the context length and $s$ is a learned scale parameter. This multiplier has the effect of sharpening/focusing the attention distribution. li2025information adopt a similar approach: controlling the entropy of the attention distribution for better length generalization.
@@
Efficient Long-Context Training and Inference: A key observation of xiong2023effective is that continual pretraining on long contexts, after initial pretraining on shorter sequences, is often sufficient and much more computationally efficient. Thus, a simple strategy for improving long-context performance, given sufficient computational budget, is continual pretraining on longer contexts. This has been adopted widely (grattafiori2024llama; gao2024train; lieber2024jamba; yang2024qwen2; cohere2025command; llama4team2025; liu2024deepseek). Of course, this strategy considerably increases the complexity and memory cost of the pretraining pipeline, making approaches like ours that can zero-shot generalise very valuable. Furthermore, one might expect long-context training to be far easier (e.g. requiring fewer long-context training steps for optimal performance) for approaches that already have good long-context performance due to zero-shot generalisation.
For inference on sequences exceeding the trained context length, alternative strategies bypass applying dense attention over the entire context, allowing for ‘infinite attention’ (munkhdalai2024leave; liu2024reattention; martins2021infty; chen2025edgeinfinite; ding2023longnet). These include maintaining a fixed-size attention window (beltagy2020longformer), retrieving relevant context tokens using heuristics like top-K proximity (han2023lm), or vector similarity search akin to episodic memory systems (fountas2024human; xiao2024infllm). While effective, these approaches operate at a higher level, managing context rather than modifying the core attention mechanism’s ability to process it directly, which is the focus of our work.
@@ 3 Methods
Following FlexAttention (dong2024flex), we define the attention “score” as the dot product of a query $q$ and keys $K$,
@@ 3.1 Formal definitions of scale-invariant total attention and attention sparsity
The $\Theta(1)$” is big-$\Theta$ notation, and means there exist constants $c_{1},c_{2}>0$ and $t_{0}$ such that for all $t>t_{0}$, $c_{1}\leq\operatorname{E}\left[Z_{t}^{t\Delta}\right]\leq c_{2}$.
@@
Figure 2: Expected entropy of scale-invariant attention at different scales is sub-logarithmic. Here, we sample sequences of independent standard Gaussian logits, and apply the scale-invariant attention transformation. We estimate the expected entropy in ranges $t,t\Delta)$, where the size of the range is controlled by $t$ (x-axis) and $\Delta$ (line color). We see that this expected entropy measure scales sub-logarithmically (left), and with the right plot suggesting a $\sim\sqrt{\log(t)}$ scaling. The dashed lines show a best linear fit.
@@ 3.2 What characteristics of the logits are required for scale-invariant attention?
then we have scale-invariant unnormalised attention sparsity (Def. 3.2). To construct an attention mechanism that satisfies Eq. (12) and Eq. (13), we consider a simplified setting in which the logits are marginally Gaussian and arise from taking Gaussian “base logits”, $\bar{L}_{t}\sim\mathcal{N}\left(0,1\right)$, and transforming them by multiplying by $a_{t}$ and adding a bias, $m_{t}$,
@@
For this solution to be valid, we only require $\beta\geq\alpha\log\alpha$ since $\log(t/\tau+1)\geq 0$ when $t\geq 0$. We formally summarise the above results in Theorem 16 (proof in Appendix F), which tells us that this approach does indeed give scale-invariant total attention and scale-invariant unnormalised attention sparsity. {restatable}theoremthmscaleinv Suppose attention logits $\{L_{t}\}$ are marginally Gaussian with mean $m_{t}$ and standard deviation $a_{t}$ defined by Eq. (16). Assuming $\alpha,\beta,\tau>0$, $\beta\geq\alpha\log\alpha$, then we have scale-invariant total attention and scale-invariant unnormalised attention sparsity. We therefore propose to scale the logits in real attention using $a_{t}$ and $m_{t}$ defined in Eq. (16). As $t$ increases, the variance, $a_{t}^{2}$, increases as the logarithm of $t$, while the mean decreases as the logarithm of $t$.
@@ 4 Experiments
In this section, we compare scale-invariant attention with other dense attention methods including Dynamic NTK interpolation (RoPE+NTK) (bloc97_2023), LogN scaling/SSMax (nakanishi2025scalable), $p$-RoPE (barbero2024round), and ALiBi (press2021train). Our results show that our method, scale-invariant $p$-RoPE, has uniformly lower validation loss at a variety of training lengths (4k, 16k, 64k). Additionally, scale-invariant $p$-RoPE demonstrates stronger zero-shot long-context generalization (e.g. in ‘Train @4k/Val @16k’ and ‘Train @4k/Val @64k’ settings) versus all other methods. Finally, scale-invariant $p$-RoPE, along with LogN, saturated“needle-in-a-haystack” task.
We pretrained GPT-2-style models (radford2019language) (with QK-norm, ReLU2 activations, etc. (modded_nanogpt_2024)) from scratch on the FineWeb dataset (penedo2024fineweb), using a fixed training data ordering and a 10M token validation set. We trained linear layers with Muon (jordan2024muon), and remaining parameters with Adam. We implemented scale-invariant attention using FlexAttention (dong2024flex).
We trained two models: one with 162M parameters and another with 304M. The 162M one were trained for 4578 steps on 2.4B tokens over a range of context lengths (4k, 16k, 64k). The 304M parameter models were trained only on the best length-generalising methods, for 10.9k steps on 10B tokens, at 4k context length. We give further experimental details in Appendix G. For the 162M parameter model, we targeted a validation loss of 3.28, following Karpathy’s GPT-2 reproduction (karpathy_llmc_481; modded_nanogpt_2024), shown in the figures as a horizontal grey dashed line.
@@
To assess whether long-context information retrieval capabilities suffered, we fine-tuned models on a needle-in-a-haystack task (note that fine-tuning on this task is unusual, but we found that prompting alone was not sufficient to perform this task, as bigger models/more pretraining would be required). Needle-in-a-haystack (kamradt2023needle) measures a model’s ability to precisely retrieve specific details (needles) from a large body of text (the haystack). Our tasks constructs prompts by concatenating text samples from the C4 dataset (roberts2019exploring) and embedding “needles” of the form ‘The special magic <city> number is <7_digit_number>’. We insert three (rather than one) needles uniformly, at random, into each context for more signal per example, with each needle contributing separately to the overall accuracy. A successful retrieval requires the model to output both the city and the associated number correctly. We trained models on sequences of length 4k, and tested at 4k, 16k, and 64k.
@@
Table 1: Final mean validation losses ($\pm$1 standard error across 3 seeds) for different methods when training with 4k context length on a 162M parameter GPT-2-style model. The error bars are small due to consistent training data ordering, and a fixed validation set.
@@
RoPE 0.962 $\pm$ 0.003 0.000 $\pm$ 0.000 0.000 $\pm$ 0.000
pp-RoPE 0.966 $\pm$ 0.001 0.250 $\pm$ 0.020 0.000 $\pm$ 0.000
NoPE 0.964 $\pm$ 0.000 0.303 $\pm$ 0.239 0.000 $\pm$ 0.000
RoPE+NTK 0.962 $\pm$ 0.001 0.217 $\pm$ 0.078 0.000 $\pm$ 0.000
YaRN 0.969 $\pm$ 0.001 0.000 $\pm$ 0.000 0.000 $\pm$ 0.000
@@ 5 Limitations
In this work, we evaluated our methods by pretraining with 162M and 304M parameter models, and investigated 7B parameter models via continual pretraining in Appendix I.5. Ideally, we would have pretrained from scratch at the multi-billion parameter scale, in line with contemporary state-of-the-art LLMs. While compute resource constraints prevented this, our results, together with the natural theoretical approach, offer no indication that our conclusions would fail to generalise to larger, commercially deployed models.
@@ 6 Conclusions
7 Acknowledgements
We sincerely thank the Engineering and Physical Sciences Research Council and the COMPASS CDT for funding Ben Anson at the University of Bristol. Many of our experiments used computational resources at the Advanced Computing Research Centre at the University of Bristol, with GPUs generously funded by Dr. Stewart.
@@ Appendix A Relating the Indicies in the Paper to Standard Indices
where $\lambda$ indexes the feature, $i$ indexes the query (i.e. the token we’re generating now) and $j$ indexes the key (i.e. the token we’re attending to). We fix $i$, and take t=ijt=i-j, so
@@ Appendix B $\operatorname{E}\left[X^{k}\exp(\alpha X)\right]$ where $X$ is Gaussian
$\mathbb{E}[X^{k}e^{\alpha X}]$ $=\int_{-\infty}^{\infty}x^{k}e^{\alpha x}\frac{1}{\sqrt{2\pi\sigma^{2}}}e^{-\frac{(x-\mu)^{2}}{2\sigma^{2}}}dx$ (23)
$=\int_{-\infty}^{\infty}x^{k}\frac{1}{\sqrt{2\pi\sigma^{2}}}e^{\alpha x-\frac{(x-\mu)^{2}}{2\sigma^{2}}}dx$ (24)
$=\int_{-\infty}^{\infty}x^{k}\frac{1}{\sqrt{2\pi\sigma^{2}}}e^{-\frac{1}{2\sigma^{2}}(x^{2}-2\mu x+\mu^{2}-2\alpha\sigma^{2}x)}dx$ (25)
$=\int_{-\infty}^{\infty}x^{k}\frac{1}{\sqrt{2\pi\sigma^{2}}}e^{-\frac{1}{2\sigma^{2}}(x^{2}-2(\mu+\alpha\sigma^{2})x+\mu^{2})}dx$ (26)
$=e^{\frac{\alpha^{2}\sigma^{2}}{2}+\alpha\mu}\int_{-\infty}^{\infty}x^{k}\frac{1}{\sqrt{2\pi\sigma^{2}}}e^{-\frac{(x-(\mu+\alpha\sigma^{2}))^{2}}{2\sigma^{2}}}dx$ (27)
@@ Appendix C Deriving $a_{t}$ and $m_{t}$ in the Logit Transformation
$\text{and }a_{t}^{2}$ $=\frac{\beta}{\alpha}-m_{t}.$ (35)
@@ G.1 Pretraining from scratch
Our base model is a modded-nanogpt (modded_nanogpt_2024) variant, which is similar to GPT-2 (radford2019language), but has the following main differences from GPT-2: RMSNorm for layer normalization (applied before the attention and MLP blocks, as well as to the token embeddings and the final output layer), squared ReLU activations, and QK Normalization (RMSNorm on query and key projections). All models used a vocabulary size of 50304, with text tokenized with the GPT-2 tokenizer (wolf2020transformers). We implemented scale-invariant attention and ALiBi using FlexAttention (dong2024flex).
@@
[…]
@@
Dataset. We used subsets of FineWeb (penedo2024fineweb) for pretraining from scratch: a 10B token subset for the 162M model, and a 100B token subset for the 304M model (from which $\sim$10B tokens were used for its training). We keep the training data ordering fixed, and we keep the validation set identical across all runs to reduce variance.